5. Solving the following using Laplace transforms: (a) y – 6y + 9y = 4te, y(0) = 1, y(0) = 0 (b) y + 2y + y = 8(t – 1),I have the answers for these I just need the detail solution on how to solve this
Answer #1

ه - (a) ,اء اور تاع امه = ۹۶ + اونه - (6) transform leplere Apply ا ا ا ا ا عر ب او يا - 4 ا ا ا ه ا 1 و 4 إ ا - إ ا . - (5(5) trylty=8Lt), y(o)=0, 4/2012 Apply Inplace transton Lh9+ 2y + 76 = Ihsct-119 Lfs(t-1)} thyhet al fylų+ L496= 5 yes) - 54